Question 13 2 pts A research report states “t(8) = 2.00, p > .05.” For this test, r2 = 2/10. O False O True
Answer #1

Solution:

Given:

t(8) = 2.00

thus

t =2.00

df = 8

then

pt = 1

72 = 72 +8

= 4-8

r^{2}=\frac{4}{12}

F | پتے

thus this is not equal to 2/10

thus this is False.